# Comparing Styles across Languages

### v1

@@ 1 Introduction
Stylistically appropriate language is crucial to successful communication within a certain culture. However, multilingual LMs sometimes struggle to generate language that is stylistically appropriate in non-English languages Hershcovich et al. (2022); Ersoy et al. (2023); Zhang et al. (2022). Standard training methods for multilingual models lead to little stylistic variation in generated text across languages Pires et al. (2019); Libovický et al. (2020); Muller et al. (2021), and multilingual systems rarely address these socially-driven factors of language Hovy and Yang (2021). As a result, downstream applications of these systems, like chatbots, are not as usable or beneficial to a non-American audience Bawa et al. (2020).
@@
Figure 1 shows an example comparison of English and Chinese politeness using this framework. We explain differences in politeness using both lexical categories and dialogue acts, with Figure 1 highlighting three chosen lexical categories. We make all code and data available publicly. [^3] https://anonymous.4open.science/r/multilingual_politeness-4302/
@@ 2.2 Feature Set Aggregation
Token-to-word grouping
@@
Category-level importances
@@ 3.1 Data Collection & Annotation Overview
Annotation
To label the dataset, we use Prolific to source annotators. Respondents are required to be native speakers of the language they annotate, as well as indicate that it is their primary spoken language day-to-day. We include attention and fluency checks to ensure our annotations are of high quality. Annotators use a 5-point scale for labeling: “Rude”, “Slightly Rude”, “Neutral”, “Slightly Polite”, and “Polite”, with three annotators labeling each utterance. We observe an average annotator agreement (Fleiss’ kappa) of 0.186 across languages. Given the highly subjective nature of politeness, we expect to see a score in this range. Additionally, we convert the Stanford Politeness Corpus Danescu-Niculescu-Mizil et al. (2013) to a 5-point scale and observe a Fleiss’ kappa of 0.153, indicating our agreement aligns with past work.
A key distinction from the Stanford Politeness Corpus and TYDIP is that we do not normalize the scores of each annotator to be centered at neutral. Levels of politeness vary culturally Leech (2007); Pishghadam and Navari (2012); we therefore do not make any assumptions about the politeness level of an average utterance. Figure A6 provides a visualization of the final annotated dataset. We observe that the average utterance from English, Spanish, and Chinese is closest to neutral, while the average Japanese utterance is closer to slightly polite.
@@ 4 Comparing Politeness via PoliteLex
When purifying our expanded lexica in Spanish and Japanese, we use the full set of 41,000 scraped talk pages to remove uncorrelated words from each category. Next, we fine-tune XLM-RoBERTa models Conneau et al. (2020) on our holistic politeness dataset (see Appendix A.4 for training details) and use the SHAP library Lundberg and Lee (2017) to extract Shapley values for each utterance. We then use Feature Set Aggregation to calculate importance scores for each PoliteLex category.
@@
Results
Table 1 analyzes our generated politeness lexica in Spanish and Japanese. We measure lexical coverage for all utterances in our dataset, and in both cases, the lexica generated by MLC has better coverage than 1:1 machine translation.
@@ Results
For certain categories, we see a strong similarity across languages. Apologetic expressions (e.g. ‘‘I’m sorry’’, ‘‘my bad’’) and expressions of gratitude (e.g. ‘‘thank you’’, ‘‘I appreciate it’’) tend to universally indicate politeness, with Japanese speakers expressing gratitude and apologies with the highest frequently.
@@ 6 Related Work
Explaining Style
One line of work builds on existing techniques Lundberg and Lee (2017); Ribeiro et al. (2016) to explain style within a single LM Aubakirova and Bansal (2016); Wang et al. (2021). Another line of work interprets style LMs by comparing learned features to those humans would consider important Hayati et al. (2021), mapping feature attributions to topics Havaldar et al. (2023), and training models to output relevant features alongside their predictions Hayati et al. (2023).
Politeness

### v2

@@ 1 Introduction
Stylistically appropriate language is crucial to successful communication within a certain culture. However, multilingual LMs sometimes struggle to generate language that is stylistically appropriate in non-English languages Hershcovich et al. (2022); Ersoy et al. (2023); Zhang et al. (2022). Standard training methods for multilingual models lead to little stylistic variation in generated text across languages Pires et al. (2019); Libovický et al. (2020); Muller et al. (2021), and multilingual systems rarely address these socially-driven factors of language Hovy and Yang (2021). As a result, downstream applications of these systems, like chatbots, are not as usable or beneficial to a non-American audience Bawa et al. (2020); Havaldar et al. (2023a).
@@
Figure 1 shows an example comparison of English and Chinese politeness using this framework. We explain differences in politeness using both lexical categories and dialogue acts, with Figure 1 highlighting three chosen lexical categories. We make all code and data available publicly. [^3] https://github.com/shreyahavaldar/multilingual_politeness
@@ 2.2 Feature Set Aggregation
Token-to-word grouping.
@@
Category-level importances.
@@ 3.1 Data Collection & Annotation Overview
Annotation.
To label the dataset, we use Prolific to source annotators. Respondents are required to be native speakers of the language they annotate, as well as indicate that it is their primary spoken language day-to-day. We include attention and fluency checks to ensure our annotations are of high quality. Annotators use a 5-point scale for labeling: “Rude”, “Slightly Rude”, “Neutral”, “Slightly Polite”, and “Polite”, with three annotators labeling each utterance. We observe an average annotator agreement (Fleiss’ kappa) of 0.186 across languages.
Given the highly subjective nature of politeness, we expect to see a score in this range. Additionally, we convert the Stanford Politeness Corpus Danescu-Niculescu-Mizil et al. (2013) to a 5-point scale and observe a Fleiss’ kappa of 0.153, indicating our agreement aligns with past work.
A key distinction between our dataset and the Stanford Politeness Corpus/TYDIPis that we do not normalize the scores of each annotator to be centered at neutral. Levels of politeness vary culturally Leech (2007); Pishghadam and Navari (2012); we therefore do not make any assumptions about the politeness level of an average utterance.
Figure A6 provides a visualization of score distributions in our final annotated dataset. We observe that the average utterance from English, Spanish, and Chinese is closest to neutral, while the average Japanese utterance is closer to slightly polite.
@@ 4 Comparing Politeness via PoliteLex
When purifying our expanded lexica in Spanish and Japanese, we use the full set of 41,000 scraped talk pages to calculate internal correlation and remove uncorrelated words from each category.
Next, we fine-tune XLM-RoBERTa models Conneau et al. (2020) on our holistic politeness dataset (see Appendix A.4 for training details) and use the SHAP library Lundberg and Lee (2017) to extract Shapley values for each utterance. Finally, we apply Feature Set Aggregation to calculate importance scores for each PoliteLex category.
@@
Dataset coverage.
Table 1 analyzes our generated politeness lexica in Spanish and Japanese. We measure dataset coverage for each language – an utterance is “covered” if it contains at least one word in a lexical category, and we define dataset coverage as the percent of covered utterances. In both cases, the lexica generated by MLC has better coverage than 1:1 machine translation using Google Translate.
Results.
@@
For certain categories, we see a strong similarity across languages. Apologetic expressions (e.g. ‘‘I’m sorry’’, ‘‘my bad’’) and expressions of gratitude (e.g. ‘‘thank you’’, ‘‘I appreciate it’’) tend to universally indicate politeness. Interestingly, we see Japanese speakers using expressions of gratitude and apology with the highest frequency across languages.
@@
6 Ablation Analysis
The PoliteLex category-level importances in Figure 4 and dialogue act importances in Figure 5 are dependent on the Shapley values extracted from fine-tuned XLM-RoBERTa models. In this section, we analyze the effect of using alternate models and training paradigms.
Effect of model size.
To investigate the role of LM size and architecture, we fine-tune Llama-2-7b Touvron et al. (2023) to analyze politeness. Comparing the results from Llama-2-7b to the results from XLM-RoBERTa, we notice stability in the direction of importance score (i.e. positive, negative, and neutral lexical categories/dialog acts are stable across both LMs). Interestingly, we observe differences in the magnitude of importance score (e.g. Llama-2-7b sees the ”Greeting” category in Chinese to be a larger indicator of politeness than XLM-RoBERTa does).
Effect of language-specific pretraining.
To investigate the role of language-specific pretraining, we fully fine-tune four RoBERTa models trained on only their respective languages. Similar to Llama-2-7b, we observe high similarity in the direction of the importance score compared to XLM-RoBERTa. We also notice much more similarity in the magnitude of importance scores. This may be due to inherent similarities between all RoBERTa models (parameter size, training methods, training data, etc.), which do not exist between XLM-RoBERTa and Llama-2-7b.
Our ablation analysis reveals that different language models pay attention to the same markers when learning to predict politeness, but learn to weigh these markers in different ways. Overall, we notice stability in which lexical categories and dialog acts indicate politeness vs. rudeness. This suggests that our groupings for feature set aggregation are both stable and successful. Section C contains additional details.
@@ 7 Related Work
Explaining style.
One line of work builds on existing techniques Lundberg and Lee (2017); Ribeiro et al. (2016) to explain style within a single LM Aubakirova and Bansal (2016); Wang et al. (2021). Another line of work interprets style LMs by comparing learned features to those humans would consider important Hayati et al. (2021), mapping feature attributions to topics Havaldar et al. (2023b), and training models to output relevant features alongside their predictions Hayati et al. (2023).
Politeness.
@@ A.4 Model Fine-Tuning
Setting Model Name
Monolingual English roberta-base Liu et al. (2019)
Monolingual Spanish bertin-roberta-base-spanish De la Rosa et al. (2022)
Monolingual Chinese chinese-roberta-wwm-ext Cui et al. (2020)
Monolingual Japanese japanese-roberta-base Cho and Sawada (2021)
Multilingual Llama-2-7b Touvron et al. (2023)
Table A5: LMs used in our ablation studies. We experiment with a different multilingual model to investigate the role of model size and architecture, as well as four separate monolingual models to investigate the role of language-specific pretraining.
@@ Appendix B Comparing Politeness Across Languages: Full Results
Appendix C Ablation Analysis: Additional Details
To fine-tune Llama-2-7b with the computational resources available to us, we use parameter efficient fine-tuning Liu et al. (2022). Specifically, we use the PEFT library Mangrulkar et al. (2022) to fine-tune a subset of the weights. We use a learning rate of $51e-$ 5 and default LORA parameters. We achieve similar performance to the metrics shown in Table A3.
To fine-tune the monolingual models (shown in Table A5), we use an identical set-up as Table A3, as the architecture of monolingual RoBERTa models and multilingual XLM-RoBERTa models is identical. Again, we observe similar performance to the metrics shown in Table A3.
We provide the resulting heatmaps in our repository: https://github.com/shreyahavaldar/multilingual_politeness